OpenStudy (anonymous):

what number must be added to complete the square x^2-1x

jhonyy9 (jhonyy9):

so how you think -1 from what numbers you will can getting ? than just make theyr product so and this product you need adding for get the cpmplete square

OpenStudy (anonymous):

The pattern used for this is \[(x+a)^2=x^2+2ax+a^2\]

jhonyy9 (jhonyy9):

complete - sorry

OpenStudy (anonymous):

So, we have to make what we have look like the right hand side of the above equation. Note that your b value is -1. Following the pattern, the b value is twice what we need to square for the constant term. That means we take the negative one, divide by two, and square it. When we add that to the binomial we started with we get\[x^2-x=(x^2-x+1/4)-1/4=(x-1/2)^2-1/4\]

OpenStudy (anonymous):

Note that the minus one fourth is an added term to make the equation true. In many cases, we would add to the other side of the equation instead of having a plus and minus of the same term on the same side.